OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, –2) and is parallel to the line y = –4x + 1?

Parth (parthkohli):

You know the point and the slope, so the point-slope form:\[y - y_0 = m(x - x_0)\]

OpenStudy (anonymous):

Parallel means Slopes are equal ie m1 = m2

Parth (parthkohli):

Now the points you have are \((x_0, y_0)\) and \(m\) is the slope.

OpenStudy (anonymous):

y = mx + c m = slope

Parth (parthkohli):

As @Yahoo! said, the slope will be equal.\[y - y_0 = -4(x - x_0)\]Now your point.

Parth (parthkohli):

Hello?

OpenStudy (anonymous):

Oooook so what your saying is the equation would be .. y-2=m(x-3) ??

OpenStudy (anonymous):

Im sorry guys, im lost with this stuff..

OpenStudy (whpalmer4):

Okay, I'll start from the top: parallel lines have the same slope. We know the slope of the other line (y = -4x + 1) to be m = -4, because it is in slope-intercept form (y = mx + b). So, our line will also have slope m = -4. Because we know a point our line passes through, and we know the slope, point-slope form is the most convenient. Our known point is (x0, y0) = (3, -2) and point-slope form for a line is \[y-y_0 = m(x-x_0)\] so plugging in what we know, we get \[y-(-2) = -4(x-3)\]\[y+2 = -4x +12\rightarrow y = -4x + 10\]Now, the problem asks for us to put the formula in standard form, which is Ax + By = C, with neither A nor B = 0. So, we add -4x to both sides giving us x and y terms on the left and a constant on the right, which is standard form. What is your answer?

OpenStudy (whpalmer4):

Sorry, we add 4x, not -4x!

OpenStudy (anonymous):

Ok so i almost got it, Except for... the problem asks for us to put the formula in standard form, which is Ax + By = C, with neither A nor B = 0. So, we add -4x to both sides giving us x and y terms on the left and a constant on the right, which is standard form. I dont understand the Ax+By=C part..

OpenStudy (whpalmer4):

It means you need to have x and y on the left side of the equals sign, and a number on the right. A and B are constants, and they can't equal 0.

OpenStudy (whpalmer4):

C is a constant, too. It can equal 0. Except in this problem, it doesn't :-)

OpenStudy (whpalmer4):

\[y = -4x + 10\]We need both x and y on the left side, so we add 4x to each side: \[y + 4x = -4x + 10 + 4x\]\[y +4x = 10\]Voilà, our equation in standard form!

OpenStudy (whpalmer4):

Here's a diagram showing the original line (in blue) and our new, parallel line passing through (3, -2) in purple.

OpenStudy (anonymous):

Ok so im still kinda confused lol im so sorry, math is like greek to me..

OpenStudy (anonymous):

OH! the graph is very helpful, thank you

OpenStudy (whpalmer4):

Millions of Greek kids learn Greek, you can do it too :-)

OpenStudy (anonymous):

Lol.. i think i should conquer algebra 2 before greek.. lol

OpenStudy (whpalmer4):

This problem is a nice one because it involves many of the tools in the toolbox, giving you a chance to see where each one is best used, except for standard form. Standard form is convenient if you want to quickly graph a line, because it is easy to determine the x and y-intercepts from standard form, and once you know them, you plot those two points, slap down your ruler and draw a line through them. In standard form: Ax + By = C x-intercept is the point where the line crosses the x-axis, so y = 0, giving us Ax = C, so the x-intercept is at (C/A, 0). Similarly, for the y-intercept, that's the point where the line crosses the y-axis, at x = 0, so By = C, and the y-intercept is at (0, C/B). I trust you can see how I arrived at those results? For the first case, Ax + B(0) = C, Ax = C, x = C/A; for the second case, A(0) + By = C, By = C, y = C/B.

OpenStudy (whpalmer4):

Point-slope form is handy when you know a point and the slope. Slope-intercept form is handy when you know the slope and the y-intercept: if you think about it, that's just a special case of point-slope form where x0 = 0: \[y-y_0 = m(x-x_0)\] \[y = m(x - 0) + y_0\]\[y = mx +y_0\]But y0 is just the y-value of the y-intercept in this case, so y_0 = b, giving us our old friend y=mx+b.

OpenStudy (whpalmer4):

Much of this isn't really difficult, it just takes the right explanation for you to understand it, and then plenty of practice to make it all second-nature. Unfortunately, not every teacher can provide the right explanation for every student, and many students aren't willing to do the practice.

OpenStudy (anonymous):

Yes. i undersand it! u should be a teacher!!